OpenStudy (anonymous):

Aki bicycle Design has determined that when x hundred bicycle are built, the average cost per bicycle by C(x)=0.2^2- 1.5x+ 7.063 where C(x) is in hundred of dollars. How many bicycle should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

The other solution: C(x)=0.2x^2-1.5x+7.063=0.2(x^2-7.5x)+7.063=0.2((x-3.5)^2 -3.5^2)+7.063 this has minimum where x-3.5=0 so x=3.5

OpenStudy (anonymous):

The 3.5 should be 3.75 I believe. Tell me if I'm wrong. Confirm by graphing the original function and finding the minimum.